I already had this question answered and webwork says the answers for the 4th and 5th question are incorrect. This is their work and I need help with this one
1 point) Susan tinds an allen antfact in the desert, where there are temperature variations from a lowin the 30s at night to a high in the 100s n the day. She is interested in how the aritact will (hotly pursued by the matary police who patrol Area 51), and slicks it in an ovent-that its, a closed bex whose lemperature she can contror precisely Let T t)be the lemperature ot the anifact Newtons law of cooling says that Tt) changes at a rate proporional to the difesence between the temperature of the environment and the temperature of the artifact This says that there is a constant k, not dependent on time, such that T= k(E-T).where Esthetemperature ofthe enmonmentme Before collecting the artact from the desert, Susan measured its temperature at a couple of tmes, and she has determined tuat tor the alen ant,ct, k = 0.55 Susan preheats her oven to 70 degrees Fahrenhet (she has stubtomy vetused to pn the metoc wons, A1 metOe oven is at exacty 70 degrees and is heating up, and the oven nuns through a tempesature cycle every 2t minutes, n wch ts temperature vaes by 20 derees above and 20degees below 70 degrees Let E(t) be the temperature of the oven ater t minutes Ad me t 0, when the antact is at a temperature of 40 degnees, she puts t n the oven Let Tt) be the temperature of the arttact at ume t Then T()-1 40 degrees Wite a dinerential equation which models the temperatune of the artfact 0 55(70- Note: Use Trather than Tt) snce the er contuses the computer Dont enter unts or ths equaton Type here to search の cs
Wrne a dfterential equation which models he temperature of the artract T = f(t,T) = Note: Use T rather than T(t)since the lather confuses the computer Dont enter unts for this equation Solve the aierential equation. To do this, you may nd it helphuilto knw that a ts a constant, then 17 18 n 15 0 55(70+20sin(t)-T) in))+C a2 +1 Aher Susan purts in the antiact in the oven, the military police break in and lake her away Think about what happens to her anract ast+o and tila in the tolowing sentence For large values oft,even though the oven temperature varies between 50 and 90 degrees, the antact vanies from 60 4593 7 9 5409 Note: You can eam partial credit on this probie You have anempted this probiem 6 tmes You have uniemted anempts remaining
ut e. 419 So 1k) yanu 바а 10土g.huni ss 4593℃ and 19 6101
Results for this submission Answer Preview 70+ 20sin(t) 40 0.55(70+ 20sin(t) - T) Result correct correct Entered 40 0.55170+20sin+T T0:443 6oy0+464991 sin t)-84453co+21.547ex 60.4593 79.6409 60.4593 79 6409 At least one of the answers above is NOT correct (1 point) Susan nds an alen artifact in the desert, where there are temperature variations from a low in the 30s at night to a high in the 100s in the day. She is interested in how the arsfact will respond to faster variations in temperature, so she kidnaps the artfact, takes it back to her lab (hoty pursued by the matary police who patrol Area 51), and stcks it in an toven mat is, a closed box whose temperature she can control precirsely. Let T(t)be the temperature of the artfact. Newoms law of cooling says mat T(t) changes at a rate proportional to the difflenence between the emperature of the environment and the temperature of the artifact. This says that there is a constant T k(E-T), where Es me temperature of the environment Ohe over) not dependent on time, such nat efore collecting the anfact trom the desert, Susan measured its bemperature at a coupie of times, and she has determined that for the asen Susan preheats her oven to 70 degrees Fahrenhet (she has stubbomly selused to join he menc word) At time Utne oven is at exactly 7U degrees and is heating up, and he oven nuns through a temperature cycle every 2t minutes, in which its temperature vanes by 20 degrees above and 20 Gegrees below 70 degrees Let E(f) be the sempecature of the oven ater f mndes
Answer #1

Solution:

We have the differential equation,

dT 0.55 (70+20sint- T) dt

dT 0.55T-0.55 (70-20 sin t) dt

The above is a first order linear differential equation we get whose integrating factor is,

0.55dt055t

multiplying the integrating factor on both sides we get,

>e0 55t(at +0.557-0.55 (70-20 sin t)

dT e0.5St0 dt + 0.55e0.55t T 0.55 (70 +20sint) e0.56

d (Te0.3 t ) dt 20 sin t) e,0.55t 0.55 (70 =

separating variables we get,

> d (Te0.st) = 0.55 (70+ 20 sint) e.Dtdt

0.55t dt d (Te0.55t) 0.55 (70+20 sin t)є

0.55t 10e zf (242 sin (t) 440 cos (t)3647) >Te 521

10 (242 sin (t)-140 cos (t)+3647) 521 ,-0.55t >T=

when t=0,T=40

=>40=dfrac{10left(-440+3647 ight)}{521}+C

> 40 61.5547 C

=>C=-21.5547

therefore,

=>T =dfrac{10left(242sinleft(t ight)-440cosleft(t ight)+3647 ight)}{521}-21.5547e^{-0.55t}

=>T =4.645sinleft(t ight)-8.445cosleft(t ight)+70-21.5547e^{-0.55t}

=>T =70+sqrt{4.645^2+(-8.445)^2}left ( sinleft(t ight)rac{4.645}{sqrt{4.645^2+(-8.445)^2}}-cosleft(t ight)rac{8.445}{sqrt{4.645^2+(-8.445)^2}} ight )-21.5547e^{-0.55t}

=>T =70+9.6381left ( sinleft(t ight)rac{4.645}{9.6381}-cosleft(t ight)rac{8.445}{9.6381} ight )-21.5547e^{-0.55t}

=>T =70+9.6381sinleft(t- an^{-1}left ( rac{8.445}{4.645} ight ) ight)-21.5547e^{-0.55t}

=>T =70+9.6381sinleft(t-61.19^o ight)-21.5547e^{-0.55t}

when t tends to infinity,

=>T_{infty} =lim_{t o infty}70+9.6381sinleft(t-61.19^o ight)-21.5547e^{-0.55t}

> Tx = 70 9.6381 sin (t-61.19の

=>T_{infty}=70pm 9.381

Therefore, T infinity lies in the range,

=>T_{infty}epsilon left (60.619,79.381 ight )